OpenStudy (anonymous):

the ratio of the volumes of two right circular cones of the same height and the base radii in the ratio 5:7 is

OpenStudy (anonymous):

Vol of cone is 1/3 the volume of cynder it is made of

OpenStudy (anonymous):

so, (1/3)(pi*r^2*h) \[ {V_1\over V_2}={r_1^2\over r_2^2} \]

OpenStudy (anonymous):

(r1)^2/(r2)^2=5/7

OpenStudy (anonymous):

25/49 is the answer

OpenStudy (anonymous):

right